OpenStudy (anonymous):

The time, in seconds, that it takes a pendulum to swing back and forth is modeled by the equation below. , where l is the length of the pendulum in feet What is the length of a pendulum that takes seconds to swing back and forth? F(l)=2\pi \sqrt{\frac{l}{32}} where l is the length of the pendulum in feet what is the length of a pendulum that takes 2.4\pi seconds to swing back and forth ? A)1.72 ft B)3.05 ft C)38.40 D)46.08

OpenStudy (mathstudent55):

\(F\left(l\right)=2\pi \sqrt{\frac{l}{32}}\) \(F(l) = 2.4\pi\) Set the two right sides equal and solve for \(l\).

OpenStudy (anonymous):

still dont understand

OpenStudy (mathstudent55):

Let's call the time t. Your equation then is \(t =2\pi \sqrt{\frac{l}{32}}\) You are asked to find the length, \(l\), for a given time of \(2.4\pi \) seconds. Since you equation is solved for \(t\), time, let \(t\) be \(2.4\pi \), and solve for \(l\). \(2.4\pi =2\pi \sqrt{\frac{l}{32}}\)

OpenStudy (mathstudent55):

You need to solve the last equation above for \(l\).

OpenStudy (mathstudent55):

Start by switching sides to get \(l\) on the left side: \(2\pi \sqrt{\frac{l}{32}}= 2.4\pi \) The next step is to divide both sides by \(2\pi\).

OpenStudy (anonymous):

let me try

OpenStudy (anonymous):

the answer would be d 46.08 right

OpenStudy (mathstudent55):

Correct

OpenStudy (mathstudent55):

\(2\pi \sqrt{\frac{l}{32}}= 2.4\pi\) \(\dfrac{\cancel{2\pi}}{\cancel{2 \pi}} \sqrt{\frac{l}{32}}= \dfrac{1.2~\cancel{2.4\pi}}{\cancel{2 \pi}}\) \(\sqrt{\frac{l}{32}}= 1.2\) \((\sqrt{\frac{l}{32}})^2= (1.2)^2\) \(\dfrac{l}{32} = 1.44\) \(l = 46.08\)